OpenStudy (anonymous):

Solve log x = 4 i got 40 is that correct

OpenStudy (anonymous):

@JA1

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

What base for the log?

OpenStudy (anonymous):

10

terenzreignz (terenzreignz):

Then raise 10 to both sides... \[\Large 10^{\log x }= x = \color{red}?\]

OpenStudy (anonymous):

so will it be x^10?

terenzreignz (terenzreignz):

Here's an example : \[\Large \log_2 x = 5\] It's base 2, so we raise 2 to both sides... \[\Large 2^{\log_2x}=2^5\]\[\Large x = 2^5\]\[\Large \color{blue}{x = 32}\]

OpenStudy (anonymous):

oh ok i see thanks

OpenStudy (anonymous):

@terenzreignz I got 10,000 for my answer!!!

terenzreignz (terenzreignz):

That's correct :P

OpenStudy (anonymous):

thanks for your help